The mean weight of 800 college students is 56 kg and the standard deviation is 5 kg. Assuming that the weight is normally distributed, determine how many students weigh:



1
Expert's answer
2021-11-03T11:59:13-0400

The mean weight of 800 college students is 56 kg and the standard deviation is 5 kg. Assuming that the weight is normally distributed, determine how many students weigh: a. Between 60 kg and 70 kg. b. More than 68 kg.

"\mu=56 \\

\sigma= 5"

a.

"P(60<X<70) = P(X<70) -P(X<60) \\

= P(Z < \frac{70-56}{5}) -P(Z < \frac{60-56}{5}) \\

= P(Z< 2.8) -P(Z< 0.8) \\

= 0.9974 -0.7881 \\

= 0.2093 \\

N = 0.2093 \times 800 = 167"

b.

"P(X> 68) = 1 -P(X< 68) \\

= 1 -P(Z< \frac{68 -56}{5}) \\

= 1 -P(Z< 2.4) \\

= 1 -0.9918 \\

= 0.0082 \\

N = 0.0082 \times 800 = 6.56 ≈ 7"